OpenStudy (anonymous):

The track of a gramophone record is the shape of a spiral curve and may be considered as a number of concentric circles of inner and outer radius 5.25cm and 10.5cm respectively. The record rotates at 33 and 1/3 revs/min and takes 18 minutes to play. Find the length of the track.

OpenStudy (kropot72):

Circumference of track C is found from the following:\[C=\pi d\]where d is the diameter. The average diameter of tracks d(av) is found as follows: \[d(av)=\frac{5.25\times 2+10.5\times 2}{2}=\frac{31.5}{2}cm\] The average circumference C(av) is as follows: \[C(av)=\pi \times \frac{31.5}{2}\] Total number of revolutions to play the record R(total) is found from the following: \[R(total)=33\frac{1}{3}\times 18=\frac{100\times 18}{3}\] The total length of the track Len(total) is found by multiplying the average circumference of one track by the toal number of revolutions as follows: \[Len(total)=\frac{100\times 18\times \pi \times31.5}{3\times 2}=?cm\]

OpenStudy (anonymous):

thanks